Use the following information to answer the question. Round your answers if necessary, to two decimal places.

Firm A can acquire firm B for $78,750 in cash or with stock worth $78,750 priced at its current price of $25 per share of stock. The synergy value of the deal is $15,000. Both firms are 100% equity financed.

Firm A: Number of Shares = 10,000 ; Price per Share = $25.00

Firm B: Number of Shares = 10,000 ; Price per Share = $10.00

What is the NPV of acquiring firm A when stock financing is used?
Answer #1

NPV = (market value of firm B + synergy benefits) - cost of aquisition

cost of aquisition = 78,750

Market value of B = 10,000*10 = 100,000

synergy benefits = 15,000

NPV = (100,000 + 15,000) - 78750

= $36,250